# Distributed Model Checking in Graphs Classes

### v1

@@ 6.2 No Extension to First Order Logic on Graphs of Bounded Degeneracy
The result for CONGEST algorithms and graph of degeneracy $2$ was already shown in Korhonen et al. [56], which in turn use a construction of Druker et al. [15]. Roughly, the lower-bound is based on a reduction to $\textsf{Disj}_{m}$, for $m=\Theta(n^{1+1/2})$. Each instance $A,B\in[m]$ is associated with a $O(n)$-node graph $G_{A,B}$ of degeneracy $2$. The set of vertices of $G_{A,B}$ are divided in two sets of vertices, denoted $V_{A}$ (and called the nodes owned by Alice) and $V_{B}$ (the nodes owned by Bob), satisfying the following conditions:
@@ 7 Conclusion
In this paper, we have established that, for every FO formula $\varphi$, and every graph class $\mathcal{G}$ of bounded expansion, there is a distributed model checking algorithm deciding whether $G\models\varphi$ in $\mathcal{O}(D+\log n)$ rounds in the CONGEST model, for every graph $G\in\mathcal{G}$. In the distributed framework, this result can be viewed as an analog of the FPT algorithms for model-checking FO formulas on classes of bounded expansion derived in [16] for centralized computing. A natural question is whether our result could be extended to a more general class of sparse graphs. In the centralized model checking, the fixed-parameter tractability of FO was extended from graphs of bounded expansion to a more general, namely nowhere dense, class of graphs in [49]. Note that nowhere dense graphs represent the ultimate limit of the program in designing fixed-parameter tractable algorithms for model checking FO formulas. Indeed, as proven in [16], for any graph class $\mathcal{G}$ that is not nowhere dense (i.e., $\mathcal{G}$ is somewhere dense), and is closed under taking subgraphs, model checking FO formulas on $\mathcal{G}$ is not fixed-parameter tractable (unless $\text{FPT}=\text{W}[1]$). Where the tractability border for CONGEST is unknown. The question is therefore twofold: (1) Is it possible to design model checking algorithms for FO formulas running in $\tilde{O}(D)$ rounds in CONGEST for nowhere dense graphs of diameter $D$? (2) Is the class of nowhere dense graphs the ultimate limit of this program under the CONGEST model?
@@
There are several difficulties in extending our approach to nowhere dense graphs. First, it is unclear how to implement the combinatorial arguments from [49], particularly the constructions of sparse neighborhood covers, in distributed model. Second, the proof of [49] builds on a weak variant of Gaifman local form and not a quantifier elimination procedure. While this is sufficient for fixed-parameter tractability of FO model-checking, we do not know whether it could be adapted in the distributed setting.
As far as our specific results are concerned, two questions arise. First, could we remove the additional logarithmic term from the complexity of our model checking algorithm? This term reflects the coloring algorithm’s complexity in [68] for graph classes of bounded expansion. So, the question is whether it is possible to compute such a coloring in $\mathcal{O}(D)$ rounds, or even in a constant number of rounds. Another intriguing question is already asked in [68], which we report here. The question is whether it is possible, for every graph class of bounded expansion, to design an $\mathcal{O}(\log n)$-round algorithm in CONGEST for identifying all nodes satisfying a “local” FO formula $\varphi$, i.e., to mark all nodes $v$ such that $\varphi(v)$ holds, where $\varphi(x)$ refers only to nodes at constant distance from $x$.

### v2

@@ 7.2 No Extension to First Order Logic on Graphs of Bounded Degeneracy
The result for CONGEST algorithms and graph of degeneracy $2$ was already shown in Korhonen et al. [47], which in turn use a construction of Druker et al. [17]. Roughly, the lower-bound is based on a reduction to $\textsf{Disj}_{m}$, for $m=\Theta(n^{1+1/2})$. Each instance $A,B\in[m]$ is associated with a $\mathcal{O}(n)$-node graph $G_{A,B}$ of degeneracy $2$. The set of vertices of $G_{A,B}$ are divided in two sets of vertices, denoted $V_{A}$ (and called the nodes owned by Alice) and $V_{B}$ (the nodes owned by Bob), satisfying the following conditions:
@@ 8 Conclusion
In centralized model checking, the fixed-parameter tractability of FO was extended from graphs of bounded expansion to the more general class of nowhere dense graphs in [39]. Notably, nowhere dense graphs represent the ultimate frontier in the program of designing fixed-parameter tractable algorithms for model checking FO formulas. Indeed, as shown in [18], for any graph class $\mathcal{G}$ that is not nowhere dense (i.e., $\mathcal{G}$ is somewhere dense) and closed under taking subgraphs, FO model checking on $\mathcal{G}$ is not fixed-parameter tractable (unless $\text{FPT}=\text{W}[1]$). The tractability frontier for FO model checking in the CONGEST model remains unknown. Moreover, there is no reason to expect that the tractability landscape in the distributed setting should mirror the one of the centralized setting. This raises several concrete and intriguing questions, discussed hereafter.
@@
Is it possible to design model-checking algorithms for FO formulas running in $\tilde{O}(D)$ rounds in CONGEST on nowhere dense graphs of diameter $D$? From a theoretical perspective, a negative answer to this question would be even more insightful than a positive one. More broadly, what is the precise boundary for $\tilde{O}(D)$-round FO model checking in the CONGEST model? Does it coincide with the class of bounded expansion graphs, extend to nowhere dense graphs, or even include a larger family of graphs? There are several difficulties in extending our approach to nowhere dense graphs. First, it is unclear how to implement the combinatorial arguments from [39], particularly the constructions of sparse neighborhood covers, in a distributed model. Second, the proof of [39] builds on a weak variant of Gaifman’s local form, and not a quantifier elimination procedure. While this is sufficient for fixed-parameter tractability of FO model-checking, we do not know whether it could be adapted in the distributed setting.
Another question is whether a meta-theorem with $\tilde{O}(D)$ round complexity can be obtained for some extensions of FO logic on planar graphs. Theorem 7.2 shows that we cannot hope to extend Theorem 1.2 to MSO, even on planar graphs. On the other hand, in centralized computing, the fixed-parameter tractability of logics lying strictly between FO and MSO is an active area of research. For example, the extension of FO known as separator logic, denoted FOL+conn, was introduced independently by Schirrmacher, Siebertz, and Vigny [60], and by Bojańczyk and Pilipczuk [3]. This logic extends FO, for every $k\geq 1$, with the general predicate $\mathsf{conn}_{k}(x,y,z_{1},\ldots,z_{k})$, which evaluates to true on a graph $G$ if the nodes corresponding to $x$ and $y$ are connected by a path that avoids the nodes corresponding to $\{z_{1},\ldots,z_{k}\}$. Another example is disjoint paths logic, denoted FOL+DP, which extends FO with the atomic predicate $\mathsf{dp}_{k}(x_{1},y_{1},\ldots,x_{k},y_{k})$, expressing the existence of $k$ pairwise internally vertex-disjoint paths between $x_{i}$ and $y_{i}$ for $i\in\{1,\ldots,k\}$. The main obstacle to developing distributed meta-theorems for these logics lies in the absence of distributed counterparts of the core algorithmic tools used in the centralized setting—such as tree decompositions which underlies tractability results for FOL+conn [58], or the irrelevant vertex technique used for FOL+DP [35].